OpenStudy (codingking03):

MEDAL:If the moon was made of gold would it be heavier than the earth?(formulas and steps pls no direct answers)

random231 (random231):

heavier? In science and engineering, the weight of an object is usually taken to be the force on the object due to gravity. Now the space is a place where gravity is zero. So how are you going to find the weight of the moon even if you know its mass.

OpenStudy (owen3):

you have to find the mass of a moon made of gold. do you know the volume of the moon?

OpenStudy (codingking03):

i have the volume of the moon (21.9 billion cubic kilometers) or (2190000 cubic centimeters)

OpenStudy (owen3):

Then use the formula mass = density * volume

OpenStudy (codingking03):

okaay one moment

OpenStudy (owen3):

In other words $$ \rm mass ~of~ gold~ moon= density~ of~ gold * volume~ of~ moon$$

OpenStudy (codingking03):

42310800 cm ^3

OpenStudy (codingking03):

and the earth is: 1.08 x 1012

OpenStudy (owen3):

We need to have same units to make a good comparison.

OpenStudy (codingking03):

soz i ment 1.08 x 1012^12 cubic km soo 1.09296 × 1018^18 cubic cm

OpenStudy (owen3):

That looks right, I will check on paper. Give me one minute.

OpenStudy (codingking03):

ok

OpenStudy (owen3):

I got a different answer.

OpenStudy (owen3):

21.9 billion cubic kilometers = 21.9 * 10^9 km^3

OpenStudy (codingking03):

what is that in cm?

OpenStudy (owen3):

density of gold = 19.32 g /cm^3 21.9 billion cubic kilometers = 21.9 * (10^9) (km^3) = 2.19 × 10^25 cm^3 mass of gold moon = 19.32 g/cm^3 × 2.19 × 10^25 cm^3

OpenStudy (codingking03):

hmm im not seeing this i used a calculator for my calculations can you give me a step by step of how yoiu got this?

OpenStudy (owen3):

Do you agree that $$ \large \rm 21.9 ~billion ~km^3= 21.9 \times 10^9 ~km^3 $$

OpenStudy (owen3):

since 1 thousand = 1 x 10^3 1 million = 1 x 10^6 1 billion = 1 x 10^9 1 trillion = 1 x 10^12

OpenStudy (codingking03):

okay so i got no it would not is that what you have?

OpenStudy (owen3):

\[\large \rm 21.9 ~billion ~km^3 \\ \large = 21.9 \times 10^9 ~km^3 \\ \large = 21.9 \times 10^9 ~ ( 1 \times 10^3 ~m ) ^3 \\ \large = 21.9 \times 10^9 ~ ( 1 \times 10^3 \times 10^2 ~cm ) ^3 \\ \large = 21.9 \times 10^9 ~ ( 10^5 cm ) ^3 \\ \large = 21.9 \times 10^9 ~ 10^{5\times 3} ~cm^3 \\ \large = 21.9 \times 10^9 ~ 10^{15}~ cm^3 \\ \large = 21.9 \times 10^{24} ~cm^3 \\ \large = 2.19 \times 10^{25} ~cm^3 \] multiply this by the density of gold, 19.32 g/cm^3 4.23*10^26 g is the mass of the gold moon the mass of earth is 5.27 x 10^27 g so earth has more mass than a gold moon

OpenStudy (owen3):

And you are correct - just read what you typed.

OpenStudy (codingking03):

ok thx

OpenStudy (owen3):

@random231 makes an interesting point

OpenStudy (owen3):

The question would be better to ask which object has more mass. Not which is heavier, since you can't weigh earth on earth. However you can imagine weighing earth and the solid gold moon on a planet much bigger than earth or the gold moon. The earth will come out 'heavier' on such a planet. \[ \rm Weight = mass \times gravity ~constant \]